OpenStudy (anonymous):

Determine the maximum and minimum values of f(x)=2sinx+sin(2x) on the interval 0,2pi

OpenStudy (campbell_st):

well the derivative is easy \[f'(x) = 2\cos(x) + 2\cos(2x) \] set f'(x) and solve for x in the given domain I'd let \[\cos(2x) =2\cos^2(x) - 1\] so it means you need to solve \[4\cos^2(x) + 2\cos(x) - 2 = 0\] you have a quadratic equation that can be factorised after dividing each term by 2 \[2\cos^2(x) + \cos(x) -1 = 0\]

OpenStudy (campbell_st):

now you should be able to find the stationary points, or critical points. You can test the points to determine which is a max and which is a min by using the 2nd derivative test. hope this helps

OpenStudy (anonymous):

Thanks for the help! This is the only problem that seemed like a blur to me. How can I split this up to find the critical points?

OpenStudy (anonymous):

I factored the expression but I do not know what values to test

zepdrix (zepdrix):

Oh you got it factored? Ok good. So were you able to determine the critical points for x? Having trouble coming up with test points?

OpenStudy (anonymous):

I have (2cos(x)-1)(cos(x)+1) and then I got cosx=1/2 and cosx=-1 and I am stuck here

zepdrix (zepdrix):

Let's look at the second term a moment. cos x = -1 Think back to your unit circle, when is this true? What angle? :O

zepdrix (zepdrix):